# On the geometric phase and its role in the design of elastic topological materials

### v1

@@ 3.1 Topological geometric phase: 1D waveguide with adiabatically-varying cross section
The perturbations $(\Delta x,\Delta y)$ define a two-dimensional parameter space illustrated in Fig. 1b. The origin represents an equilateral triangular cross section. A uniform waveguide with this cross section has fully degenerate flexural modes that in the long wavelength approximation correspond to the in-plane and the out-of-plane bending modes. This point in parameter space (and its corresponding values) is a full degeneracy. Tracing out the loop $L$ in the parameter space of Fig. 1b leads to the waveguide in Fig. 1c whose cross section at longitudinal coordinate $z$ (in the physical space) resembles Fig. 1a.
Figure 1: A topological geometric phase is observed in 1D waveguides with adiabatically varying cross sections along the longitudinal $z$ direction. (a) The shaded region is the cross section of the waveguide at a given $z$. The vertices $\mathbf{r}_{1}$, $\mathbf{r}_{2}$, and $\mathbf{r}_{3}$ form an equilateral triangle. $\mathbf{R}$ perturbs the cross section from an equilateral triangle. (b) The parameter space is comprised of the components of $\mathbf{R}$, namely, $\Delta x$ and $\Delta y$. The origin of the parameter space corresponds to an equilateral triangle, which is a full degeneracy. (c) A waveguide generated by tracing loop $L$ in the parameter space. The cross sections and points corresponding on the loop are marked 1-4. (d) A loop in the parameter space encircling the full degeneracy. (e,f) A front view and an orthographic view of the waveguide corresponding to (d). (g) A loop in the parameter space not encircling the full degeneracy. (h,i) A front view and an orthographic view of the waveguide corresponding to (g). Flexural guided modes are excited at the left end of the waveguides in (f) and (i), and the response is recorded at the right end. The longitudinal $z$ coordinate has been scaled down by a factor of $60$ in (f) and (i) for better visualization. (j) The $y$ component of the displacement field response at a point on the end of the waveguide is plotted for both waveguides.
@@
Results show that the two waveguides produce time series that are perfectly out-of-phase with each other (Fig. 1j). This phase difference of $\pi$ is due to the geometric phase acquired by the waveguide of Fig. 1f. (The choice of parameter variation causes both waveguides to accumulate the same amount of dynamical phase, so the difference in response is purely geometrical.)
@@
An additional novel example of a topological geometric phase in shear waves propagating in an anisotropic bulk is presented in Sec. 1 of the supplementary material.
@@ 3.2 Nontopological geometric phase: Polarized waves in helical waveguides
This section explores an example of elastic waveguides that gives rise to a geometric phase of nontopological nature. Consider waveguides supporting a pair of fully degenerate flexural guided modes that have orthonormal displacement amplitudes $\mathbf{U}^{(1)}(x,y)$ (Fig. 2a) and $\mathbf{U}^{(2)}(x,y)$ (Fig. 2b), which correspond to horizontal and vertical particle displacements respectively. Orthonormality implies $\int_{S}\mathbf{U}^{(i)*}\cdot\mathbf{U}^{(j)}\,\mathrm{d}x\mathrm{d}y=\delta_{ij}$, where $S$ is the cross section, $\delta_{ij}$ is the Kronecker delta, $i=1,2$, and $j=1,2$. A linear combination of these displacement amplitudes, $\mathbf{U}=c_{1}\mathbf{U}^{(1)}+c_{2}\mathbf{U}^{(2)}$, represents a guided mode with the particle displacements along the unit vector $\hat{\mathbf{c}}=(c_{1}\hat{\mathbf{e}}_{x}+c_{2}\hat{\mathbf{e}}_{y})/\sqrt{c_{1}^{2}+c_{2}^{2}}$. (Fig. 2c). The plane wave solution corresponding to $\mathbf{U}$ is typically referred to as a polarized wave. The unit vector $\hat{\mathbf{c}}$ is called the polarization vector (Fig. 2d). The angle subtended between $\hat{\mathbf{c}}$ and the $c_{1}$ axis is the polarization angle $\gamma$ of the wave. A wave with displacement amplitude proportional to $\mathbf{U}^{(1)}$, shown in Fig. 2a, ($\mathbf{U}^{(2)}$, Fig. 2b) is said to be horizontally (vertically) polarized as its polarization vector marked by the red (blue) arrow is horizontal (vertical) in Fig. 2d.
@@
Figure 2: A nontopological geometric phase manifests in polarized waves propagating in helical waveguides. The amplitude of displacement function $\mathbf{U}$ of flexural guided modes in a uniform waveguide with circular cross section is plotted in (a)-(c). The out-of-plane component of displacement is negligible. (a) A horizontally polarized guided mode. (b) A vertically polarized guided mode. (c) A guided mode with a general polarization. (d) Polarization vectors for (a) in red, (b) in blue, and (c) in green. The polarization angle of (c) is marked as $\gamma$. (e) A path $C$ on the spherical parameter space parameterized by $s$. 00, $s_{0}$, and $l$ denote the path coordinate $s$ at the marked locations. (f) Path $P$ (dashed black line) created by the curve $C$ of (e) and the waveguide generated by sweeping a circular cross section along $P$. The unit tangent $\hat{\mathbf{t}}(s_{0})$ of path $P$ at a path coordinate $s_{0}$ corresponds to the point on the sphere $\hat{\mathbf{t}}(s_{0})$ on curve $C$ in (e). (g) Loops $L_{1}$ and $L_{2}$ in the parameter space. $\theta$ is the half-cone angle subtended by loop $L_{1}$. (h) A helical waveguide generated from curve $L_{1}$. A horizontally polarized flexural wave is excited at the left end. As it propagates down the waveguide, it acquires a polarization angle $\gamma$. The radius has been scaled up by a factor of 10 for better visualization.
@@
Polarized waves provide several examples of nontopological geometric phases [42, 39, 40], where the underlying physics resembles the motion of precession in a Foucault pendulum [40, 19]. Additional examples of nontopological geometric phases are provided by surface waves [36, 37, 38] and waves in rotating media [41, 37].
@@
We illustrate the nontopological geometric phase by means of helical waveguides created by sweeping a circular cross section through a generating path $P$ (in real space) [42, 39]. Such waveguides support polarized flexural waves as each cross section normal to the generating path is circular and, thus, it supports fully degenerate flexural modes. The geometry of a waveguide illustrated in Fig. 2f is specified by the radius $r$ of the cross section (fixed to be 0.05 $m$), the length $l$ of the centerline of the waveguide, and a curve $C$ on a spherical parameter space in Fig. 2e. $l$ determines the length of the generating path $P$ in Fig. 2f. Curve $C$ of Fig. 2e specifies the unit tangent vector $\hat{\mathbf{t}}(s_{0})$ of the generating path $P$ at the path coordinate $s_{0}$ in Fig. 2f. Together, $l$ and $C_{1}$ specify the generating path $P_{1}$ along which the circular cross section of radius $r$ is swept to generate the waveguide in Fig. 2f. $l$ is assumed large enough to ensure adiabatic variation of the path tangents. In what follows, the curve on the unit sphere is assumed to be a closed loop, such as curve $L_{1}$ in Fig. 2g. This assumption implies that the initial and final tangent vectors of the generating path are parallel. Then, as a wave propagates along the waveguide, the direction of propagation given by $\hat{\mathbf{t}}$ adiabatically completes a cycle.
@@
As in the previous section, we can show the occurrence and properties of the geometric phase in helical waveguides by exploiting numerical simulations. Consider the helical waveguide depicted in Fig. 2h generated by the loop $L_{1}$ in Fig. 2g. The generating helical path has a radius $r_{h}=$ 2.4 $m$, pitch $p=$ 14 $m$, and consists of one revolution. The length of the path is $l=\sqrt{(2\pi r_{h})^{2}+p^{2}}=$ 20.58 $m$. The angle between the unit tangent vector and the axis of the helix denoted by $\theta$ is constant (Fig. 2g). It equals $\theta=\cos^{-1}(p/l)=47.12^{\circ}$ [2]. The reference displacement amplitude functions $\mathbf{U}^{(1)}$ and $\mathbf{U}^{(2)}$ of two fully degenerate and orthonormal guided modes corresponding to horizontal and vertical polarization are shown in Figs. 2a,b. One end of the waveguide is excited harmonically to produce a horizontally polarized flexural wave. The displacement field response is recorded at the other end of the waveguide. The recorded displacement amplitude function is plotted in Fig. 2c, indicating that the polarization angle does not return to zero at the end of the waveguide. The polarization angle is computed as $\gamma=114^{\circ}$ by projecting the displacement profile onto the horizontally and vertically polarized guided modes (Fig. 2d). This polarization angle is a geometric phase. As will be discussed in Sec. 4.2, the geometric phase is analytically computed as the solid angle subtended by the loop $L_{1}$. The analytical result is $\gamma_{\mathrm{theory}}=-2\pi\cos\theta\mod 2\pi=-244.9^{\circ}\mod 360^{\circ}=115.1^{\circ}$, which agrees well with the numerical computation.
@@
By repeating the simulation with a greater value of generating path length $l$, the geometric phase is verified to be independent from the rate of variation of the parameters (as long as the adiabatic condition is satisfied). However, the geometric phase is not topologically protected. Repeating the simulation after perturbing the curve $L_{1}$ to $L_{2}$ changes the polarization angle to $129^{\circ}$. (See Sec. 3 of the supplementary material for the simulations.) Indeed, changing the loop in the spherical parameter space changes the solid angle it subtends, which in turn changes the geometric phase. This is why the geometric phase discussed here is classified as “nontopological”.
@@ 3.3 Summary of the general characteristics of the geometric phase
The defining characteristics of the geometric phase in elastic waveguides are summarized as follows. First, waves in elastic waveguides that adiabatically vary along the direction of wave propagation can exhibit geometric phases either as a correction to the dynamical phase or to the polarization angle. Second, a geometric phase depends solely on the path described by the parameters in the parameter space, and it is independent of the rate of variation of these parameters. This property justifies the terminology “geometric”. Third, geometric phases are categorized into two classes: topological and nontopological. Topological geometric phases are associated with full degeneracies in an eigenvalue problem (Eqs. (8), (9)). Physically, if a waveguide exhibits a topological geometric phase, the value of the geometric phase will not be affected by small perturbations to either the geometry or the composition (e.g. the material distribution) of the waveguide. On the other hand, if a waveguide exhibits a nontopological geometric phase, the value of the geometric phase will change with perturbations to the properties of the waveguide.
@@ 4.1 Fiber bundles
Figure 3: (a) A two-dimensional manifold $\mathcal{M}$. At a point $\mathbf{R}$, a 1D fiber $\mathcal{F}$ can be attached to create the fiber bundle in (c). (b) A point $\mathbf{R}$ of the manifold $\mathcal{M}$ is in a neighborhood $\mathcal{U}$ that is equivalent to the Euclidean plane $\mathbb{R}^{2}$. (c) A fiber bundle $\mathcal{E}$ created from a two-dimensional base manifold $\mathcal{M}$ and 1D fiber $\mathcal{F}$. The pictured fiber bundle is trivial. The marked point $u$ has a base coordinate $\mathbf{R}$ and fiber coordinate $f$. (d) $u$ lies in the neighborhood $\mathcal{U}\times\mathcal{F}$. (e) A sphere with a tangent vector $\mathbf{v}$ attached at the point $\mathbf{R}$. The green plane is a tangent plane attached at $\mathbf{R}$. Attaching the tangent plane at each point gives rise to the tangent bundle of a sphere $TS^{2}$.
@@
The physical relevance of fiber bundles lies in modeling the response of position- or parameter-dependent systems. The space of allowable values of positions or parameters gives rise to the base manifold. At each position or parameter, the space of possible responses of the system is modeled as a fiber. For example, the volume of an elastic body defines a manifold $\mathcal{M}$. Each point of the elastic body is associated with a displacement vector, which can be modeled as an element of the fiber $\mathbb{R}^{3}$. A point $u$ on the resulting fiber bundle has a base coordinate $\mathbf{R}$, which describes a point on the elastic body, and a fiber coordinate $\mathbf{u}$, which describes the displacement at $\mathbf{R}$.

### v3

@@ 3.1 Topological geometric phase: 1D waveguide with adiabatically-varying cross section
The perturbations $(\Delta x,\Delta y)$ define a two-dimensional parameter space illustrated in Fig. 1b. The origin represents an equilateral triangular cross section. A uniform waveguide with this cross section has fully degenerate flexural modes that, in the long wavelength approximation, correspond to the in-plane and the out-of-plane bending modes. This point in parameter space (and its corresponding values) is a full degeneracy resulting from symmetry. Tracing out the loop $L$ in the parameter space of Fig. 1b leads to the waveguide in Fig. 1c whose cross section at the longitudinal coordinate $z$ (in the physical space) resembles Fig. 1a.
Figure 1: A topological geometric phase is observed in 1D waveguides with adiabatically varying cross sections along the longitudinal $z$ direction. (a) The shaded region is the cross section of the waveguide at a given $z$. The vertices $\mathbf{r}_{1}$, $\mathbf{r}_{2}$, and $\mathbf{r}_{3}$ form an equilateral triangle. $\mathbf{R}$ perturbs the cross section from an equilateral triangle. (b) The parameter space is comprised of the components of $\mathbf{R}$, namely, $\Delta x$ and $\Delta y$. The origin of the parameter space corresponds to an equilateral triangle, which corresponds to a full degeneracy. (c) A waveguide generated by tracing loop $L$ in the parameter space. The cross sections and points corresponding to the loop are marked 1-4. (d) A loop in the parameter space encircling the full degeneracy. (e,f) A front view and an orthographic view of the waveguide corresponding to (d). (g) A loop in the parameter space not encircling the full degeneracy. (h,i) A front view and an orthographic view of the waveguide corresponding to (g). Flexural guided modes are excited at the left end of the waveguide in (f) and (i), and the response is recorded at the right end. The longitudinal $z$ coordinate has been scaled down by a factor of $60$ in (f) and (i) for better visualization. (j) The $y$ component of the displacement field response at a point on the end of the waveguide is plotted for both waveguides.
@@
Results show that the two waveguides produce time series that are perfectly out-of-phase with each other (Fig. 1j). This phase difference of $\pi$ is due to the geometric phase acquired by the waveguide of Fig. 1f. The choice of parameter variation causes both waveguides to accumulate the same amount of dynamical phase, so the difference in response is purely geometrical.
@@
An additional novel example of a topological geometric phase in shear waves propagating in an anisotropic bulk material is presented in Sec. 1 of supplementary material.
@@ 3.2 Nontopological geometric phase: Polarized waves in helical waveguides
This section explores an example of elastic waveguides that gives rise to a geometric phase of nontopological nature. Consider waveguides supporting a pair of fully degenerate flexural guided modes that have orthonormal eigenfunctions $\mathbf{U}^{(1)}(x,y)$ (Fig. 2a) and $\mathbf{U}^{(2)}(x,y)$ (Fig. 2b), which correspond to horizontal and vertical particle displacements respectively. Orthonormality implies $\int_{S}\mathbf{U}^{(i)*}\cdot\mathbf{U}^{(j)}\,\mathrm{d}x\mathrm{d}y=\delta_{ij}$, where $S$ is the cross section, $\delta_{ij}$ is the Kronecker delta, $i=1,2$, and $j=1,2$. A linear combination of these eigenfunctions, $\mathbf{U}=c_{1}\mathbf{U}^{(1)}+c_{2}\mathbf{U}^{(2)}$, represents a guided mode with the particle displacements along the unit vector $\hat{\mathbf{c}}=(c_{1}\hat{\mathbf{e}}_{x}+c_{2}\hat{\mathbf{e}}_{y})/\sqrt{c_{1}^{2}+c_{2}^{2}}$. (Fig. 2c). The plane wave solution corresponding to $\mathbf{U}$ is typically referred to as a polarized wave. The unit vector $\hat{\mathbf{c}}$ is called the polarization vector (Fig. 2d). The angle subtended between $\hat{\mathbf{c}}$ and the $c_{1}$ axis is the polarization angle $\gamma$ of the wave. A guided mode eigenfunction proportional to $\mathbf{U}^{(1)}$, shown in Fig. 2a, ($\mathbf{U}^{(2)}$, Fig. 2b) is said to be horizontally (vertically) polarized as its polarization vector marked by the red (blue) arrow is horizontal (vertical) in Fig. 2d.
@@
Figure 2: A nontopological geometric phase manifests itself in polarized waves propagating in a helical waveguide. The guided mode eigenfunction $\mathbf{U}$ of flexural guided modes in a uniform waveguide with circular cross section is plotted in (a)-(c). The out-of-plane component of the displacement is negligible. (a) A horizontally polarized ($x$-axis) guided mode. (b) A vertically polarized ($y$-axis) guided mode. (c) A guided mode with a general polarization. (d) Polarization vectors for (a) in red, (b) in blue, and (c) in green. The polarization angle of (c) is marked as $\gamma$. (e) A path $C$ on the spherical parameter space parameterized by $s$. 00, $s_{0}$, and $l$ denote the path coordinate $s$ at the marked locations. (f) Path $P$ (dashed black line) created by the curve $C$ of (e) and the waveguide generated by sweeping a circular cross section along $P$. The unit tangent $\hat{\mathbf{t}}(s_{0})$ of path $P$ at a path coordinate $s_{0}$ corresponds to the point on the sphere $\hat{\mathbf{t}}(s_{0})$ on curve $C$ in (e). (g) Loops $L_{1}$ and $L_{2}$ in the parameter space. $\theta$ is the half-cone angle subtended by loop $L_{1}$. (h) A helical waveguide generated from curve $L_{1}$. A horizontally polarized flexural wave is excited at the left end. As it propagates down the waveguide, it acquires a polarization angle $\gamma$. The radius has been scaled up by a factor of 10 for better visualization.
@@
Polarized waves provide several examples of nontopological geometric phases [50, 47, 48], which resemble the precession motion of a Foucault pendulum in terms of the associated parameter space and eigenfunctions [48, 19]. Additional examples of nontopological geometric phases are provided by surface waves [44, 45, 46] and waves in rotating media [49, 45].
@@
We illustrate the concept of nontopological geometric phase by means of helical waveguides created by sweeping a circular cross section through a generating path $P$ (in real space) [50, 47]. Such waveguides support polarized flexural waves, as each cross section normal to the generating path is circular, and circular cross sections support fully degenerate flexural modes [56]. The geometry of a waveguide (see Fig. 2f) is specified by the radius $r$ of the cross section (fixed to be 0.05 $m$), the length $l$ of the centerline of the waveguide, and a curve $C$ on a spherical parameter space (Fig. 2e). $l$ determines the length of the generating path $P$ in Fig. 2f. Curve $C$ in Fig. 2e specifies the unit tangent vector $\hat{\mathbf{t}}(s_{0})$ of the generating path $P$ at the path coordinate $s_{0}$ (Fig. 2f). Together, $l$ and $C_{1}$ specify the generating path $P_{1}$ along which the circular cross section of radius $r$ is swept to generate the waveguide in Fig. 2f. $l$ is assumed large enough to ensure the adiabatic variation of the path tangents. In what follows, the curve on the unit sphere is assumed to be a closed loop, such as curve $L_{1}$ in Fig. 2g. This assumption implies that the initial and final tangent vectors of the generating path are parallel. Then, as a wave propagates along the waveguide, the direction of propagation given by $\hat{\mathbf{t}}$ adiabatically completes a cycle.
@@
As in the previous section, we can show the occurrence and properties of the geometric phase in helical waveguides by exploiting numerical simulations. Consider the helical waveguide depicted in Fig. 2h generated by the loop $L_{1}$ in Fig. 2g. The generating helical path has a radius $r_{h}=$ 2.4 $m$, pitch $p=$ 14 $m$, and consists of one revolution. The length of the path is $l=\sqrt{(2\pi r_{h})^{2}+p^{2}}=$ 20.58 $m$. The angle between the unit tangent vector and the axis of the helix denoted by $\theta$ is constant (Fig. 2g) and equals $\theta=\cos^{-1}(p/l)=47.12^{\circ}$ [2]. The reference eigenfunctions $\mathbf{U}^{(1)}$ and $\mathbf{U}^{(2)}$ of two fully degenerate and orthonormal guided modes corresponding to horizontal and vertical polarizations are shown in Figs. 2a,b. One end of the waveguide is excited harmonically to produce a horizontally polarized flexural wave. The displacement field response is recorded at the other end of the waveguide. The recorded guided mode eigenfunction is plotted in Fig. 2c, indicating that the polarization angle does not return to zero at the end of the waveguide. The polarization angle is computed as $\gamma=114^{\circ}$ by projecting the displacement profile onto the horizontally and vertically polarized guided modes (Fig. 2d). This polarization angle is a geometric phase. As will be discussed in Sec. 4.2, the geometric phase is analytically computed as the solid angle subtended by the loop $L_{1}$. The analytical result is $\gamma_{\mathrm{theory}}=-2\pi\cos\theta\mod 2\pi=-244.9^{\circ}\mod 360^{\circ}=115.1^{\circ}$, which agrees well with the numerical computation.
@@
By repeating the simulation with a greater value of the generating path length $l$, the geometric phase is verified to be independent from the rate of variation of the parameters (as long as the adiabatic condition is satisfied). However, the geometric phase is not topologically protected. Repeating the simulation after perturbing the curve $L_{1}$ to $L_{2}$ changes the polarization angle to $129^{\circ}$. The corresponding simulations are reported in Sec. 3 of supplementary material. Indeed, changing the loop in the spherical parameter space changes the solid angle it subtends, which in turn changes the geometric phase. This is why the geometric phase discussed here is classified as “nontopological”.
@@ 3.3 Summary of the general characteristics of the geometric phase
The defining characteristics of the geometric phase in elastic waveguides are summarized as follows. First, waves in elastic waveguides that adiabatically vary along the direction of wave propagation can exhibit geometric phases either as a correction to the dynamical phase or to the polarization angle. Second, a geometric phase depends solely on the path described by the parameters in the parameter space, and it is independent of the rate of variation of these parameters. This property justifies the terminology “geometric”. Third, geometric phases are categorized into two classes: topological and nontopological. Topological geometric phases are associated with full degeneracies in the associated eigenvalue problem describing the dynamics of the system (Eqs. (8), (9)). Physically, if a waveguide exhibits a topological geometric phase, the value of the geometric phase will not be affected by small perturbations to either the geometry or the composition (e.g. the material distribution) of the waveguide. On the other hand, if a waveguide exhibits a nontopological geometric phase, the value of the geometric phase will change with perturbations to the properties of the waveguide.
@@ 4.1 Fiber bundles
Figure 3: (a) A two-dimensional manifold $\mathcal{M}$. At a point $\mathbf{R}$, a 1D fiber $\mathcal{F}$ can be attached to create the fiber bundle in (c). (b) A point $\mathbf{R}$ of the manifold $\mathcal{M}$ is in a neighborhood $\mathcal{U}$ that is equivalent to the Euclidean plane $\mathbb{R}^{2}$. (c) A fiber bundle $\mathcal{E}$ created from a two-dimensional base manifold $\mathcal{M}$ and a 1D fiber $\mathcal{F}$. The pictured fiber bundle is trivial. The marked point $u$ has a base coordinate $\mathbf{R}$ and fiber coordinate $f$. (d) $u$ lies in the neighborhood $\mathcal{U}\times\mathcal{F}$. (e) A sphere with a tangent vector $\mathbf{v}$ attached at the point $\mathbf{R}$. The green plane is a tangent plane attached at $\mathbf{R}$. Attaching the tangent plane at each point gives rise to the tangent bundle of a sphere $TS^{2}$. (f) A curved beam in undeformed (gray) and deformed (red) configurations. (g) Fiber bundle description of the curved beam modeled by the Euler-Bernoulli theory. The base manifold $\mathcal{M}$ is the centerline of the undeformed beam, and the fiber is $\mathbb{R}$ to model the transverse displacement of the beam. They lead to the (trivial) fiber bundle $\mathcal{E}=\mathcal{M}\times\mathbb{R}$. The marked point $u$ has the base coordinate $\mathbf{R}$ and fiber coordinate $\mathbf{u}$. Sections of the fiber bundle corresponding to the undeformed (black) and deformed (red) beam are marked.
@@
The physical relevance of the fiber bundles lies in their ability to capture the response of position- or parameter-dependent systems. The space of allowable values of either positions or parameters gives rise to the base manifold. At each position or parameter, the space of possible responses of the system is modeled as a fiber. An example of a position-dependent system modeled as a fiber bundle is the deformation of an elastic body (Fig. 3f,g). An undeformed elastic body defines a manifold $\mathcal{M}$ (Fig. 3f). Each point of the elastic body is associated with a displacement vector, which can be modeled as an element of a fiber. The fiber depends on the permissible displacements: for example, a general 3D problem requires $\mathbb{R}^{3}$ as the fiber, a plane strain problem requires $\mathbb{R}^{2}$ as the fiber, and approximate 1D theories, such as the Euler-Bernoulli beam theory, require $\mathbb{R}$ as the fiber. A point $u$ on the resulting fiber bundle $\mathcal{E}$ has a base coordinate $\mathbf{R}$, which describes a point on the elastic body, and a fiber coordinate $\mathbf{u}$, which describes the displacement at $\mathbf{R}$ (Fig. 3f,g). A deformed configuration of the body specifies $\mathbf{u}(\mathbf{R})$, which corresponds to a section of the fiber bundle. Parameter-dependent systems modeled as fiber bundles form the basis to understand the geometric phase (Sec. 5).